OpenStudy (josee):

A lighthouse is located on a small island, 4 kms away from the nearest point P on a straight shoreline, and its light makes six rev- olutions per minute. How fast is the beam of light moving along the shoreline when it is 3 kms from P ?

OpenStudy (amistre64):



OpenStudy (amistre64):

da/dt = 12pi per minute, i can see that much at least

OpenStudy (amistre64):

dP/dt = dP/da * da/dt

OpenStudy (amistre64):

What is the relation of the angle to the distance ?

OpenStudy (josee):

huh?

OpenStudy (amistre64):

tan(a)=P/4 right?

OpenStudy (josee):

Yeppp

OpenStudy (josee):

and do the tangent inverse

OpenStudy (josee):

of 3/4?

OpenStudy (amistre64):

well, at that instant it is 3/4 ; but in general

OpenStudy (amistre64):

tan(a) = P/4 sec^2(a) da/dt = 1/4 dP/dt

OpenStudy (amistre64):

dP/dt = 4 sec^2(a) da/dt

OpenStudy (amistre64):

dP/dt = 4 sec^2(a) 12pi dP/dt = 44pi sec^2(a) what is the angle of tan-1(3/4) ?

OpenStudy (amistre64):

dP/dt = 44pi sec^2(tan^(-1)(3/4)) = 215.98 units per hour maybe?

OpenStudy (josee):

hmm

OpenStudy (amistre64):

prolly not tho

OpenStudy (josee):

noo. not close either

OpenStudy (amistre64):

ive never been able to get a handle on the lighthouse one fer some reason

OpenStudy (josee):

1. 4488pi kms/h 2. 4504pi kms/h 3. 449pi kms/h 4. 4500pi kms/h 5. 4501pi kms/h

OpenStudy (amistre64):

oh, my units are in minutes

OpenStudy (amistre64):

and the pi is estimated

OpenStudy (amistre64):

(275 pi)/4; per minute 60(275 pi)/4 = 4125pi per hour

OpenStudy (amistre64):

...... harumph

OpenStudy (amistre64):

4500 .... 4*12 = 48

OpenStudy (amistre64):

dP/dt = 48pi sec^2(tan^(-1)(3/4)) = 75pi /min 75*60 = 4500pi / hr

OpenStudy (josee):

omg thxxxx!